OpenStudy (anonymous):

can anyone give me hint if sec[(x+y)/(x-y)] =a prove that dy/dx=y/x should i substitute x=ycos2theta ???

OpenStudy (anonymous):

or should i subtitute x=tantheta

hartnn (hartnn):

yes! x+y)/(x-y = sec^-1 a and since 'a' is constant, the entire right side is constant and its derivative = 0 !

OpenStudy (anonymous):

oh :OOOOOO i don't need to subtitute anything :OOO

hartnn (hartnn):

you can simplify it even more let sec^-1 a = c (just for sake of simplicity) (x+y)/ (x-y) = c now do you know about componendo -dividendo ???

OpenStudy (anonymous):

no :'( forgot

hartnn (hartnn):

okk... \(\large \dfrac{a}{b}=\dfrac{c}{d} \implies \dfrac{a+b}{a-b}=\dfrac{c+d}{c-d}\) so, what about \(\large \dfrac{x+y}{x-y}=\dfrac{c}{1}\) ??

OpenStudy (anonymous):

yo got it XD saw it on the net :)

OpenStudy (anonymous):

x+y+x-y /(x+y-x+y) =2x/2y =x/y :D

hartnn (hartnn):

lol that is just \(\dfrac{x}{y}=\dfrac{c+1}{c-1}\) we have still even did not start differentiating!

hartnn (hartnn):

\(\dfrac{x}{y}=\dfrac{c+1}{c-1}=k\) k is another constant for simplicity ... now differentiate left of x/y = k using quotient rule! and since k is constant, right side = 0

hartnn (hartnn):

@kittycat01 , still here ?

hartnn (hartnn):

oh, dear! she is gone without realizing that she had not completed the solution (didn't even start differentiating!) :P

OpenStudy (loser66):

(x/y)' how? partial derivative or y respect to x? @hartnn

hartnn (hartnn):

not partial, implicit differentiation diff. both sides w.r.t x

hartnn (hartnn):

\(\huge \dfrac{y-x \dfrac{dy}{dx}}{y^2}=0\)

OpenStudy (loser66):

and the numerator =0 solve for dy/dx?

hartnn (hartnn):

yup

OpenStudy (loser66):

Thanks.

OpenStudy (anonymous):

@hartnn -lol sorry i got so excited that i closed my net and i start doing al the sums again